OpenStudy (anonymous):

solve for x: 2/5x + 2 = 10 A.) 20 B.) 24 C.) 4 D.) 8

OpenStudy (blurbendy):

2/5x + 2 = 10 2/5x = 8 2x = 40 x = 20

OpenStudy (anonymous):

hey thxs! mind if i ask u one more?

OpenStudy (blurbendy):

sure

OpenStudy (anonymous):

5/2x + 15/2 = 15 A.) -3 B.) -1.5 C.) 0 D.) 3

OpenStudy (blurbendy):

5/2x + 15/2 = 15 5/2x = 15/2 5x = 15 x = 3

OpenStudy (anonymous):

oh got it! one more? sorry haha im just starting to get these!

OpenStudy (blurbendy):

ok

OpenStudy (anonymous):

(x+16)/3 =3x

OpenStudy (anonymous):

A.) 3 B.) 5 C.) 6 D.) 10

OpenStudy (blurbendy):

(x + 16) / 3 = 3x x + 16 = 9x 16 = 8x x = 2 I think you gave the wrong answer choices.

OpenStudy (anonymous):

omg sorry lol

OpenStudy (blurbendy):

np